# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ Abstract
Estimating the Riesz representer is a central problem in debiased machine learning for causal and structural parameter estimation. Various methods for Riesz representer estimation have been proposed, including Riesz regression and covariate balancing. This study unifies these methods within a single framework. Our framework fits a Riesz representer model to the true Riesz representer under a Bregman divergence, which includes the squared loss and the Kullback–Leibler (KL) divergence as special cases. We show that the squared loss corresponds to Riesz regression, and the KL divergence corresponds to tailored loss minimization, where the dual solutions correspond to stable balancing weights and entropy balancing weights, respectively, under specific model specifications. We refer to our method as generalized Riesz regression, and we refer to the associated duality as automatic covariate balancing. Our framework also generalizes density ratio fitting under a Bregman divergence to Riesz representer estimation, and it includes various applications beyond density ratio estimation. We also provide a convergence analysis for both cases where the model class is a reproducing kernel Hilbert space (RKHS) and where it is a neural network.
@@ 1 Introduction
The Riesz representer plays a crucial role in debiased machine learning for a variety of causal and structural parameter estimation problems (Chernozhukov et al., 2022b), such as Average Treatment Effect (ATE) estimation, Average Marginal Effect (AME) estimation, Average Policy Effect (APE) estimation, and covariate shift adaptation (Shimodaira, 2000; Uehara et al., 2020; Chernozhukov et al., 2025). The Riesz representer arises from the Riesz representation theorem for the parameter functional, and it also has a close connection to semiparametric efficiency bounds (Newey, 1994). In particular, by using the Riesz representer appropriately, we can obtain semiparametric estimators that are asymptotically linear with the efficient influence function, which is also referred to as the Neyman orthogonal score (Chernozhukov et al., 2018).
@@
To address this issue, end to end approaches for Riesz representer estimation have been explored, including Riesz regression (Chernozhukov et al., 2021). In particular, in ATE estimation, entropy balancing weights (Hainmueller, 2012), stable balancing weights (Zubizarreta, 2015), and tailored loss minimization (Zhao, 2019) have been proposed. In covariate shift adaptation, direct density ratio estimation methods have been proposed (Sugiyama et al., 2012).
@@
In the following sections, we introduce our general setup (Section 2) and then propose our Riesz representer estimation method (Section 3). We refer to Riesz representer fitting under a Bregman divergence as generalized Riesz regression. We may also refer to it as Bregman–Riesz regression or direct bias correction term estimation. However, we adopt the term generalized Riesz regression because the choice of loss function is closely related to the choice of link function. For example, the duality relationship between Riesz regression and stable balancing weights holds when we use a linear link (a linear model) for the Riesz representer. In contrast, the duality relationship between tailored loss minimization and entropy balancing weights holds when we use a logistic link (a logistic model) for the Riesz representer. We make this correspondence explicit in Section 4. We refer to this property as automatic covariate balancing.
For our proposed framework, we and subsequent works provide convergence rate analyses, major applications, and several extensions. In Section 5, we show the convergence rate of the Riesz representer model to the true Riesz representer when using reproducing kernel Hilbert space (RKHS) regression and neural networks. In particular, we establish minimax optimality. In Section 6, we provide applications of our framework to ATE, AME, and APE estimation, as well as covariate shift adaptation. In Appendices G and H, we also introduce extensions of our framework that are developed in subsequent works. Kato (2025a) points out that nearest neighbor matching-based density ratio estimation proposed in Lin et al. (2023) is a special case of least-squares importance fitting (LSIF) for density ratio estimation (Kanamori et al., 2009). Moreover, since LSIF can be interpreted as Riesz regression, nearest neighbor matching-based ATE estimation can also be interpreted as ATE estimation via Riesz regression. Kato (2025c) proposes a Riesz representer estimation method based on score matching in diffusion models (Song et al., 2021).
@@ 2 Setup
We first describe our general setup, which includes various tasks, such as ATE estimation, as special cases. We denote a pair of an outcome variable $Y\in{\mathcal{Y}}$ and regressors $X\in{\mathcal{X}}$ by
$W\coloneqq(X,Y)$
be a pair of an outcome variable $Y\in{\mathcal{Y}}$ and regressors $X\in{\mathcal{X}}$, where $\mathcal{Y}$$\in{\mathbb{R}}$ and ${\mathcal{X}}\in{\mathbb{R}}^{d}$ are outcome and ($d$-dimensional) regressor spaces, respectively. Let $P_{0}$ be the distribution that generates $W$. We assume that we can observe $n$ i.i.d. copies of $W$, denoted as
@@ 2.1 Riesz representer
For simplicity, we assume that the expected functional $\gamma\mapsto{\mathbb{E}}_{P_{0}}\big[m(W,\gamma)\big]$ is linear and continuous in $\gamma$, which implies that there is a constant $C>0$ such that ${\mathbb{E}}\big[m(W,\gamma)\big]^{2}\leq C\mathbb{E}\left[\gamma(X)^{2}\right]$ holds for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. From the Riesz representation theorem, there exists a function $v_{m}$ with $\mathbb{E}\left[v_{m}(X)^{2}\right]<\infty$ such that
@@
for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. This formulation follows Chernozhukov et al. (2022b) and can be generalized to non-linear maps $\gamma\mapsto\mathbb{E}\left[m(W,\gamma)\right]$. However, we do not present this generalization because it is not our main focus, and the linear case is sufficient for presenting our results. The function $v_{m}$ is referred to as the Riesz representer in Chernozhukov et al. (2025). In ATE estimation, the Riesz representer has also been referred to as the bias-correction term or the clever covariates (van der Laan, 2006; Schuler & van der Laan, 2024).
@@ 2.2 Neyman Orthogonal Scores
$\psi(W;\eta,\theta)\coloneqq m(W,\gamma)+\alpha(X)\big(Y-\gamma(X)\big)-\theta_{0}.$
@@
Thus, orthogonality ensures that first-order errors from estimating $\eta_{0}$ do not affect the asymptotic distribution of the final estimator $\widehat{\theta}$ of $\theta_{0}$, provided cross-fitting (or a Donsker condition) and mild convergence rate conditions on $\widehat{\eta}$ hold (Chernozhukov et al., 2018, 2022b).
@@
This estimation apparoch is tradinally called bias correction or estimation equation approach (van der Vaart, 2002; Schuler & van der Laan, 2024), which are reformulated as (automatic) debiased machine learning (Chernozhukov et al., 2018, 2022b, 2024). This estimator is a generalization of the augmetend inverse probability weighting (AIPW) estimator in causal inference (Bang & Robins, 2005; Tsiatis, 2007). For simplicity, we refer to the estiamtor $\widehat{\theta}$ as the AIPW estimator.
@@ 2.3 Examples
Let the regressor $X$ be $X\coloneqq(D,Z)$, where $D\in\{1,0\}$ is a treatment indicator, and $Z\in{\mathcal{Z}}$ is a covariate vector, where $\mathcal{Z}$} denotes its support. Following the Neyman–Rubin framework, let $Y(1)\in{\mathcal{Y}}$ and $Y(0)\in{\mathcal{Y}}$ be the potential outcomes for treated and control units. In ATE estimation, using the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$, our goal is to estimate the ATE, defined as
@@
$\psi^{\text{ATE}}\left(W;\eta,\theta^{\text{ATE}}\right)\coloneqq m^{\text{ATE}}(W,\gamma)+\alpha^{\text{ATE}}(X)\Big(Y-\gamma(X)\Big)-\theta^{\text{ATE}},$
with $\eta=(\gamma,\alpha^{\text{ATE}})$. Plugging in estimates of $\eta_{0}$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{ATE}}\left(W_{i};\widehat{\eta},\widehat{\theta}^{\text{ATE}}\right)=0$, we can obtain an ATE estimator as
$).\widehat{\theta}^{\text{ATE}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\alpha_{0}^{\text{ATE}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{ATE}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
In practice, $\alpha_{0}^{\text{ATE}}$ is replaced by an estimator. This estimator is called the augmented inverse probability weighting (AIPW) estimator, or the doubly robust estimator.
@@
In our framework, we estimate $\alpha^{\text{AME}}_{0}$ and $\gamma_{0}$ using generalized Riesz regression and then construct a Neyman-targeted estimator.
@@
Let $P_{X,0}$ be the marginal distribution of $X$ under $P_{0}$. When $P_{X,0}$ is absolutely continuous with respect to $P_{X,1}$ with density ratio $r_{0}(x)\coloneqq\frac{p_{1}}{p_{0}}(x)$, the Riesz representer is
@@
The Neyman orthogonal score
$\psi^{\text{CS}}(W;\eta,\theta)=\gamma(X)+\alpha^{\text{CS}}(X)\big(Y-\gamma(X)\big)-\theta$
delivers a debiased estimator by combining source residuals with target averaging, and it accommodates independent source and target samples via cross fitting or data fusion.
@@
In causal inference, Uehara et al. (2020) investigates efficient ATE estimation and policy learning under a covariate shift. Kato et al. (2024b) applies the method for adaptive experimental design, and Kato et al. (2025) exntends the framework for PU learning setup. Sugiyama et al. (2008) consider covariate shift adaptation using importance weighting estimated by LSIF, which are equivalent to Riesz regression in density ratio estimation (Kato, 2025b). Chernozhukov et al. (2025) and Kato et al. (2024a) investigate efficient estimation of parameters under a covariate shift from some different perspectives.
@@
$(\alpha-1)^{2}/2$ LSIF Riesz regression (Riesz net and Riesz forest)
@@
$\big(|t|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ UKL divergence minimization Tailored loss minimization
@@
$\qquad-(|\alpha|+C)\log(|\alpha|+C)$ (Qin, 1998) (Standard)
@@ 3 Generalized Riesz Regression
In this study, we generalize existing methods through the lens of Bregman divergence minimization. The Bregman divergence is a general discrepancy measure that includes the squared loss and the KL divergence as special cases. In this section, we propose Riesz representer fitting under a Bregman divergence, which we also refer to as generalized Riesz regression or Bregman-Riesz regression. The term generalized Riesz regression reflects the fact that the choice of loss function is closely connected to the choice of link function from the viewpoint of covariate balancing. We explain this viewpoint in Section 4 and refer to it as automatic covariate balancing. This section provides a general formulation, and we introduce applications of generalized Riesz regression in Section 6.
@@ 3.1 Bregman Divergence
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{H}}}\widehat{\text{BD}}_{g}\big(\alpha\big)+\lambda J(\alpha),$ (1)
@@
The choice of the regularization function is important because Riesz representer estimation is known to exhibit a characteristic overfitting phenomenon, often described as train-loss hacking or the density chasm. For details, see Appendix F.

### v2

@@ Abstract
Estimating the Riesz representer is central to debiased machine learning for causal and structural parameter estimation. We propose generalized Riesz regression, a unified framework that estimates the Riesz representer by fitting a representer model via Bregman divergence minimization. This framework includes the squared loss and the Kullback–Leibler (KL) divergence as special cases: the former recovers Riesz regression, while the latter recovers tailored loss minimization. Under suitable model specifications, the dual problems correspond to covariate balancing, which we call automatic covariate balancing. Moreover, under the same specifications, outcome averages weighted by the estimated Riesz representer satisfy Neyman orthogonality even without estimating the regression function, a property we call automatic Neyman orthogonalization. This property not only reduces the estimation error of Neyman orthogonal scores but also clarifies a key distinction between debiased machine learning and targeted maximum likelihood estimation. Our framework can also be viewed as a generalization of density ratio fitting under Bregman divergences to Riesz representer estimation, and it applies beyond density ratio estimation. We provide convergence analyses for both reproducing kernel Hilbert space (RKHS) and neural network model classes. A Python package for generalized Riesz regression is available at https://github.com/MasaKat0/grr.
@@ 1 Introduction
The Riesz representer plays a crucial role in debiased machine learning for a variety of causal and structural parameter estimation problems (Chernozhukov et al., 2022b), such as Average Treatment Effect (ATE) estimation (Imbens & Rubin, 2015), Average Marginal Effect (AME) estimation, Average Policy Effect (APE) estimation, and covariate shift adaptation (Shimodaira, 2000; Uehara et al., 2020; Chernozhukov et al., 2025). The Riesz representer arises from the Riesz representation theorem for the parameter functional, and it also has a close connection to semiparametric efficiency bounds (Newey, 1994). In particular, by using the Riesz representer appropriately, we can obtain semiparametric estimators that are asymptotically linear with the efficient influence function, which is also referred to as the Neyman orthogonal score (Chernozhukov et al., 2018).
@@
To address this issue, end-to-end approaches for Riesz representer estimation have been explored, including Riesz regression (Chernozhukov et al., 2021). In particular, in ATE estimation, entropy balancing weights (Hainmueller, 2012), stable balancing weights (Zubizarreta, 2015), and tailored loss minimization (Zhao, 2019) have been proposed. In covariate shift adaptation, direct density ratio estimation methods have been proposed (Sugiyama et al., 2012).
@@
In the following sections, we introduce our general setup (Section 2) and then propose our Riesz representer estimation method (Section 3). We refer to Riesz representer fitting under a Bregman divergence as generalized Riesz regression. We may also refer to it as Bregman–Riesz regression, direct bias correction term estimation, generalized tailored loss minimization, or generalized covariate balancing (Remark Remark). However, we adopt the term generalized Riesz regression because the choice of loss function is closely related to the choice of link function. For example, the duality relationship between Riesz regression and stable balancing weights holds when we use a linear link (a linear model) for the Riesz representer. In contrast, the duality relationship between tailored loss minimization and entropy balancing weights holds when we use a logistic link (a logistic model) for the Riesz representer. We make this correspondence explicit in Section 4. We refer to this property as automatic covariate balancing.
For our proposed framework, this paper provides convergence rate analyses, major applications, and several extensions. In Section 6, we show the convergence rate of the Riesz representer model to the true Riesz representer when using reproducing kernel Hilbert space (RKHS) regression and neural networks. In particular, we establish minimax optimality. In Section 7, we provide applications of our framework to ATE, AME, and APE estimation, as well as covariate shift adaptation. In Appendices G and H, we also introduce extensions of our framework that are developed in subsequent works. Kato (2025c) points out that nearest neighbor matching-based density ratio estimation proposed in Lin et al. (2023) is a special case of least-squares importance fitting (LSIF) for density ratio estimation (Kanamori et al., 2009). Moreover, since LSIF can be interpreted as Riesz regression, nearest neighbor matching-based ATE estimation can also be interpreted as ATE estimation via Riesz regression. Kato (2025e) proposes a Riesz representer estimation method based on score matching in diffusion models (Song et al., 2021).
@@ 2 Setup
We first describe our general setup, which includes various tasks, such as ATE estimation, as special cases. We denote the pair $W\coloneqq(X,Y)$, where $Y\in{\mathcal{Y}}$ is an outcome and $X\in{\mathcal{X}}$ is a regressor vector. Here, $\mathcal{Y}$$\subseteq{\mathbb{R}}$ and ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ are outcome and ($d$-dimensional) regressor spaces, respectively. Let $P_{0}$ be the distribution that generates $W$. We assume that we can observe $n$ i.i.d. copies of $W$, denoted as
@@ 2.1 Riesz representer
For simplicity, we assume that the expected functional $\gamma\mapsto{\mathbb{E}}\big[m(W,\gamma)\big]$ is linear and continuous in $\gamma$, which implies that there is a constant $C>0$ such that ${\mathbb{E}}\big[m(W,\gamma)\big]^{2}\leq C\mathbb{E}\left[\gamma(X)^{2}\right]$ holds for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. From the Riesz representation theorem, there exists a function $v_{m}$ with $\mathbb{E}\left[v_{m}(X)^{2}\right]<\infty$ such that
@@
for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. The function $v_{m}$ is referred to as the Riesz representer in Chernozhukov et al. (2025). In ATE estimation, the Riesz representer has also been referred to as the bias-correction term or the clever covariates (van der Laan, 2006; Schuler & van der Laan, 2024).
Remark.
This formulation follows Chernozhukov et al. (2022b) and can be generalized to non-linear maps $\gamma\mapsto\mathbb{E}\left[m(W,\gamma)\right]$. However, we do not present this generalization because it is not our main focus, and the linear case is sufficient for presenting our results.
@@ 2.2 Neyman Orthogonal Scores
$\psi(W;\eta,\theta)\coloneqq m(W,\gamma)+\alpha(X)\big(Y-\gamma(X)\big)-\theta.$
@@
Thus, orthogonality ensures that first-order errors from estimating $\eta_{0}$ do not affect the asymptotic distribution of the final estimator $\widehat{\theta}$ of $\theta_{0}$, provided cross fitting (or a Donsker condition) and mild convergence rate conditions on $\widehat{\eta}$ hold (Chernozhukov et al., 2018, 2022b).
@@
This estimation approach is traditionally called one-step bias correction, or the estimating equation approach (van der Vaart, 2002; Schuler & van der Laan, 2024). Recent work reformulates it as (automatic) debiased machine learning (Chernozhukov et al., 2018, 2022b, 2022c, 2024). This estimator generalizes the augmented inverse probability weighting (AIPW) estimator in causal inference (Bang & Robins, 2005; Tsiatis, 2007).
@@ 2.3 Examples
Let the regressor $X$ be $X\coloneqq(D,Z)$, where $D\in\{1,0\}$ is a treatment indicator, and $Z\in{\mathcal{Z}}$ is a covariate vector, where $\mathcal{Z}$} denotes its support. Following the Neyman–Rubin framework, let $Y(1),Y(0)\in{\mathcal{Y}}$ be the potential outcomes for treated and control units. In ATE estimation, using the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$, our goal is to estimate the ATE, defined as
@@
$\psi^{\text{ATE}}\left(W;\eta,\theta\right)\coloneqq m^{\text{ATE}}(W,\gamma)+\alpha^{\text{ATE}}(X)\Big(Y-\gamma(X)\Big)-\theta.$
Plugging in estimates of $\eta^{\text{ATE}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{ATE}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{ATE}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an ATE estimator as
$).\widehat{\theta}^{\text{ATE}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{ATE}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{ATE}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
This estimator is called the augmented inverse probability weighting (AIPW) estimator, or the doubly robust estimator.
@@
Plugging in estimates of $\eta^{\text{AME}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{AME}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{AME}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an AME estimator as
$).\widehat{\theta}^{\text{AME}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{AME}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{AME}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
@@
Plugging in estimates of $\eta^{\text{APE}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{APE}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{APE}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an APE estimator as
$).\widehat{\theta}^{\text{APE}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{APE}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{APE}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
@@
Let $P_{X,0}$ be the marginal distribution of $X$ under $P_{0}$, and $p_{0}(x)$ and $p_{1}(x)$ be the pdfs. Assume that if $p_{0}(x)>0$, then $p_{1}(x)>0$ holds. Then, the Riesz representer is
@@
The Neyman orthogonal score is given as
$\psi^{\text{CS}}(W;\eta,\theta)=\gamma(X)+\alpha^{\text{CS}}(X)\big(Y-\gamma(X)\big)-\theta.$
Plugging in estimates of $\eta^{\text{CS}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{CS}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{CS}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an estimator as
$).\widehat{\theta}^{\text{CS}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{CS}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{CS}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
Density ratio estimation
@@
Sugiyama et al. (2008) consider covariate shift adaptation using importance weighting estimated by LSIF, which are equivalent to Riesz regression in density ratio estimation (Kato, 2025d). Chernozhukov et al. (2025) and Kato et al. (2024a) investigate efficient estimation of parameters under a covariate shift from some different perspectives.
@@
$(\alpha-C)^{2}$ LSIF Riesz regression (Riesz net and Riesz forest)
@@
$\big(|\alpha|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ UKL divergence minimization Tailored loss minimization ($\alpha=\beta=-1$)
@@
$\qquad-(|\alpha|+C)\log(|\alpha|+C)$ (Qin, 1998) (Standard approach)
Tailored loss minimization ($\alpha=\beta=0$)
(Zhao, 2019)
@@ 3 Generalized Riesz Regression
In this study, we generalize existing methods through the lens of Bregman divergence minimization. The Bregman divergence is a general discrepancy measure that includes the squared loss and the KL divergence as special cases. In this section, we propose Riesz representer fitting under a Bregman divergence, which we also refer to as generalized Riesz regression. The term generalized Riesz regression reflects the fact that the choice of loss function is closely connected to the choice of link function from the viewpoint of covariate balancing. We explain this viewpoint in Section 4 and refer to it as automatic covariate balancing. This section provides a general formulation, and we introduce applications of generalized Riesz regression in Section 7.
Remark.
We also refer to our method as direct bias correction term estimation, Bregman Riesz regression, or generalized tailored loss minimization. In an earlier draft, we used the term direct bias correction term estimation because the Riesz representer is almost equivalent to the bias correction term in one step bias correction. Bregman Riesz regression highlights that the method combines the Bregman divergence with Riesz regression. Generalized tailored loss minimization emphasizes that our generalized Riesz regression extends tailored loss minimization (Zhao, 2019) and covers a broader class of methods, including Riesz regression. As discussed in Section 5, standard covariate balancing methods implicitly assume a constant (homogeneous) ATE across $x$, whereas the covariate balancing property under generalized Riesz regression allows for heterogeneity. From this viewpoint, we call the method generalized covariate balancing.
@@ 3.1 Bregman Divergence
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g}\big(\alpha\big)+\lambda J(\alpha),$ (1)
@@
The choice of the regularization function is important because Riesz representer estimation is known to exhibit a characteristic overfitting phenomenon, often described as train-loss hacking or the density chasm. For details, see Section 8.